4a) In a Millikan oil drop apparatus, an electric field of 26 kN/C is created in a chamber to suspend oil drops that are missing electrons. What is the electric force on a drop that is missing 4 electrons? _______N

4b) The Angstrom (10-10m) is a unit that is useful in measuring distances beteen atoms. How large is the electric field at a distance of 2.4 Angstrom from a singly ionized sodium atom?

__________N/C *This would be an atom that's missing a single electron.
Answer #1

4a) E=F/q --> F=qE=4eE~166.4*10-16 N

e=1.6*10-19 C

4b) E=\frac{q}{4\pi \varepsilon _{0}d^{2}}=k\frac{q}{d^{2}}

k=\frac{1}{4\pi \varepsilon _{0}}=9\times 10^{9}\: N\cdot m^{2}/C^{2}

The charge of single ionized sodium is +e.

E~2.5*1010 V/m